# Efficient Perspective-Correct 3D Gaussian Splatting Using Hybrid Transparency

### v2

@@ 1 Introduction
However, the remarkable performance of 3DGS is achieved through a series of approximations that, while effective in many scenarios, introduce limitations in multi-view consistency. One significant issue arises from the affine approximation used in projecting 3D Gaussian splats onto the image plane, causing differing pixel contributions depending on camera placement. Another one is the per-primitive sorting, which results in patchy colors appearing and disappearing during movements (called popping [RSP24]). In order to solve the issues, it is necessary to accurately sort and evaluate Gaussians per pixel, introducing a strong bottleneck on the sorting stage. In this paper, we introduce an pipeline which improves the original 3DGS framework in eliminating these artifacts while offering superior performance through a hybrid transparency approximation.
@@
  • We evaluate our efficient implementation, which outperforms established methods significantly, while maintaining visual quality.
@@ 2.1 Novel View Synthesis
Alternatively, point-based models use explicit point clouds for geometry, rendering images through fast rasterization. Due to the discrete nature of point clouds, recent point-based radiance field methods employ large convolutional neural networks (CNN) for hole-filling in image-space [ASK20, RFS22, FRF23, KPLD21, KLR22, HFF23], but these CNNs are computationally expensive. Addressing this limitation, Franke et al. showed that assigning a radius to each point in combination with trilinear interpolation into an image pyramid allows using a smaller CNN [FRFS24]. While enabling fast rendering, image quality of point-based models is often lacking behind implicit NeRF- and grid-based models. Very recently, Hahlbohm et al. showed that point-based models and grid-based models can be combined to improve image quality and robustness [HFK24]. Remaining issues with point-based models are temporal instabilities and the reliance on dedicated GPUs, which are both due to the used CNN. Another approach involves converting a trained NeRF model into a faster representation for inference [HSM21, YLT21]. Recent methods achieve impressive frame rates with only a small quality loss via triangle mesh baking [YHR23, RGS24] or distilling of state-of-the-art NeRF models into memory-efficient, hybrid representations [DHR24]. While they render exceptionally fast even on low-end mobile devices, a major disadvantage is their reliance on the NeRF model that needs to be trained before the – often equally expensive – baking process. Other approaches that proved effective directly optimize a triangle meshes [CFHT23] or a set of tetrahedra [KS23].
@@ 2.2 Gaussian Splatting
More recently, point-based rendering regained momentum as a powerful primitive for differentiable rendering in novel view synthesis [KPLD21, ASK20, RFS22]. In 2023, Kerbl et al. [KKLD23] re-introduced rasterization-based rendering of point primitives with anisotropic Gaussian extent in their seminal work 3D Gaussian Splatting (3DGS). Utilizing fast, tile-based rasterization [LZ21], 3DGS optimizes a set of explicit 3D Gaussians via gradient descent and an adaptive density control mechanism. Due exceptionally fast rendering during inference, 3DGS has inspired a plethora of follow-up research addressing, e.g., anti-aliasing [YCH24], compression [BKL24], dynamic scenes [LKLR24, WYF24], large scenes [KMK24], and the reliance on an initial point cloud [KRS24, NMR24]. Especially relevant to this work is StopThePop by Radl et al. [RSP24], which employs a hierarchical sorting approach to reduce popping artifacts originating from approximate sorting used in 3DGS. Notably, the per-pixel ordering resulting from this hierarchical sorting approach is not guaranteed to be fully correct. We also present a method for avoiding these popping artifacts, but achieve this by changing the blending procedure to a hybrid transparency-based approach so that our method only requires partial depth-ordering of the initial $K$ splats contributing to each pixel color. Thus, we avoid the need for a sophisticated sorting implementation which, even when compared to the approximate solution used by 3DGS, improves performance. Equally important in the context of this work is the perspectively accurate 2D Gaussian Splatting (2DGS) by Huang et al. [HYC24] as well as concurrent work that introduces ray tracing into the 3DGS framework [YSG24, MLMP24]. While the latter also provide a perspectively accurate rendering formulation by using ray tracing, they introduce a significant slowdown during both the optimization and subsequent inference. Furthermore, their evaluation of the 3D Gaussian primitives relies on matrix inversion, which is prone to numerical instabilities when Gaussians fall flat along one or more principal axes. The perspectively correct approach by Huang et al. [HYC24] circumvents these numerical instabilities by utilizing an optimized approach for 2D, ie., degenerate 3D Gaussians [SWBG06, WHA07]. In this work, we present an approach that enables using perspectively correct rendering of 3D Gaussians through ray-splat intersection: we extend the perspectively correct approach of Weyrich et al. [WHA07] to work with non-degenerate 3D Gaussians, which enables the use of ray-splat intersection for rasterization-based rendering with negligible overhead and no issues with respect to numerical instabilities.
@@ 3.2 Accurate Splat Bounding and Evaluation
In 3DGS, Kerbl et al. [KKLD23] project a 3D Gaussian onto the image plane using the Jacobian of the affine approximation of the projective transformation [ZPvBG01a]. The result is a 2D Gaussian that can easily be evaluated for each pixel and also allows for easy construction of an axis-aligned bounding box around the center (e.g., at $3\sigma$) in screen space for accelerated computation. However, this approach is not fully perspective-accurate, meaning it introduces artifacts when splats are viewed from certain angles. For example, these artifacts become visible when placing the camera on top of a flat surface with the viewing direction perpendicular to the surface normal (see first row in Fig. 1).
@@
Next, we seek to compute the value of the 3D Gaussian for a given viewing ray. In computer graphics, ray intersections with deformed primitives, such as our 3D splats, are an established practice since the dawn of ray tracing, generally following the approach of inversely transforming a viewing ray into the undeformed space. In our case, re-using quantities of the bounding-box setup above, this would mean transforming the ray by $(VPMT)^{-1}$ into splat space, where the distance of the transformed ray to the origin yields $\rho$, Eq. 1. In order to avoid the matrix inversion that we successfully sidestepped during bounding-box calculation, we once again follow Weyrich et al. and represent a viewing ray through a pixel at $(x_{\text{s}},y_{\text{s}})$ as two perpendicular planes [^1] We follow the common convention of representing planes as homogeneous vectors $\boldsymbol{p}=(a,b,c,d)^{\!\top}$, so that a point $\boldsymbol{x}$ lies on the plane iff $\boldsymbol{p}^{\!\top}\boldsymbol{x}=0$. $\boldsymbol{\pi}_{x}=(1,0,0,-x_{\text{s}})^{\!\top\!}$ and $\boldsymbol{\pi}_{y}=(0,1,0,-y_{\text{s}})^{\!\top\!}$ in screen space. In contrast to 3D points, transforming those planes by $(VPMT)^{-1}$ into planes $\boldsymbol{\pi}^{\text{s}}_{x/y}$ in splat space requires the inverse-transposed mapping, that is:
@@ 4 Experiments
Zip-NeRF [BMV23] 0.828 28.56 0.219 5h 0.2 N///a 0.878 26.75 0.233 5h 0.1 N///a
3DGS [KKLD23] 0.814 27.20 0.254 18m40s 316.6 3.31M 0.866 25.27 0.276 20m18s 216.2 1.50M
StopThePop [RSP24] 0.814 27.30 0.252 12m03s 212.5 3.28M 0.866 24.99 0.270 14m27s 206.2 1.49M
2DGS [HYC24] 0.795 26.81 0.297 21m08s 134.9 2.01M 0.851 24.55 0.314 28m45s 95.8 0.95M
GOF [YSG24] 0.821 27.27 0.238 1h16m 16.7 2.91M 0.866 25.01 0.274 1h35m 12.3 1.26M
Ours 0.822 27.17 0.234 10m07s 457.3 2.19M 0.866 24.62 0.272 8m33s 380.7 0.81M
@@ 4.1 Setup
Following recent work, we evaluate on the established Mip-NeRF360 [BMV22] and Tanks and Temples [KPZK17] datasets, which provide a diverse set of 17 real scenes with varying challenges. We use all nine scenes from the Mip-NeRF360 dataset as well as all eight scenes from the intermediate set of the Tanks and Temples dataset resulting in a total of 17 scenes. This allows us to evaluate our method on a broad spectrum of challenges regarding both geometric and photometric aspects. As is common practice, we use the 4×/ 2×4$\times$/\ 2$\times$4 × / 2 × downscaled images for the outdoor///indoor scenes of the Mip-NeRF360 dataset. For the Tanks and Temples scenes, we use the full-resolution images to evaluate performance at full HD resolution. To compute meaningful quality metrics, we use the established 7:1 train///test split [BMV22] for all scenes. Central to our evaluation is the comparison against the original 3D Gaussian Splatting [KKLD23]. Furthermore, we compare against the recent StopThePop [RSP24] and 2D Gaussian Splatting [HYC24], as they also provide solutions for the problems addressed in this work. We also compare against Gaussian Opacity Fields (GOF) [YSG24], a very recent work that uses ray tracing-based volume rendering. For all baselines, we use the official implementation to compute the results for all 17 scenes and use the same script for all quality metric computations. We use a single RTX 4090 for all experiments and outline our approach for measuring frame rates in Appendix A.
@@ 4.2 Implementation and Optimization
We implement our method from scratch in PyTorch and CUDA using the implementation of Kerbl et al. [KKLD23] as a reference. While the optimization pipeline and hyperparameters are mostly the same as in 3DGS, we make multiple modifications. First, we replace the opacity reset with a decay strategy [RSP24] that multiplies splat opacity by $\lambda_{o}=0.9995$ every $50$ iterations during densification. We tackle aliasing problems by employing our own fast CUDA implementation of the 3D filter proposed by Yu et al. [YCH24]. As the screen-space gradients 3DGS uses for densification are not available with our splat evaluation scheme, we follow Moenne-Loccoz et al. and instead directly use the gradient of the splat means scaled by half the respective distance to the camera in view space [MLMP24]. Furthermore, we adapt the densification changes from GOF to our approach [YSG24] and apply visibility score-based pruning as proposed by Niemeyer et al. [NMR24]. For our blending formulation, we use a core with $K=16$ and employ a second threshold $\tau_{K}=0.05$ specifying the minimum $\alpha$ of a fragment to be considered for the core. As the insertion sort that we employ for keeping track of per-pixel core causes significant divergence, we reduce the tile size to $8\times 8$ in our implementation. Lastly, we make use of an optimized SSIM implementation [MKVC24] to speed up loss computation.
@@ 4.3 Results
The quantitative results in Tab. 1 show that our approach provides significantly faster training and rendering. Compared to the next fastest baseline, we achieve a training speedup of 1.4×1.4$\times$1.4 × and render 1.6×1.6$\times$1.6 × faster on average. In terms of image quality, we observe that our method achieves about equal results across the board. However, all 3DGS-based approaches are clearly outperformed by Zip-NeRF [BMV23]. Taking our contributions into consideration, we observe that our approach outperforms approaches with perspectively correct splat evaluation: Compared to 2DGS [HYC24], we achieve better results across the board reassuring that 3D Gaussian splats are more expressive than the 2D Gaussian surfels used by 2DGS. GOF [YSG24] achieves slightly higher image quality but requires longer training and fails to render in real-time. It should be noted that 2DGS and GOF were fine-tuned for mesh extraction after optimization, which is an interesting topic but not a focus of this work. Seeing that we achieve similar image quality but faster training and rendering compared to StopThePop [RSP24], the results confirm that our approach is a valid alternative to prevent popping artifacts for rasterization-based 3D Gaussian splat rendering.
3DGS StopThePop 2DGS Ours Ground Truth
@@
Method Preprocess Tiling Blending Total
3DGS [KKLD23] 0.451 ms 2.212 ms 2.134 ms 4.797 ms
StopThePop [RSP24] 0.649 ms 0.738 ms 3.599 ms 4.986 ms
Ours 0.556 ms 0.754 ms 2.007 ms 3.317 ms
Table 2: Breakdown of render timings at a resolution of 1920×1080 pixels measured on an RTX 4090 GPU. For our approach, we render the trained bicycle model (4.35M splats) for all test-set viewpoints and average across 100 runs. Values for 3DGS and StopThePop are copied from Radl et al. [RSP24], who use a slightly different approach for their performance benchmarks but average in a similar manner, which makes numbers comparable.
@@
3DGS 0.730 24.64 0.265 21m17s 293.2
Ours 0.742 24.62 0.237 12m09s 480.6
Ours w/o per-pixel HT 0.748 24.81 0.230 17m26s 448.7
Table 3: We showcase the potential of our perspectively correct splat evaluation by omitting our hybrid transparency-based blending in favor of the approximately depth-ordered alpha blending from 3DGS [KKLD23]. As it forgoes per-pixel ordering, this version of our model is subject to popping artifacts, and it is slightly less computationally efficient than our default configuration, but outperforms 3DGS across all metrics. The reported values are averaged results for the five outdoor scenes from the Mip-NeRF360 dataset [BMV22].
Perspectively Correct Splat Evaluation. As previously discussed, our contributions can be integrated into 3D Gaussian splatting implementations independently. Our hybrid transparency approach (Sec. 3.3) is mainly a method for addressing popping artifacts without sacrificing computational efficiency. However, our perspectively accurate splat evaluation (Sec. 3.2) should prove beneficial across a much broader range of applications due to its accuracy. As can be seen in Tab. 3, using it alongside the approximately depth-ordered alpha blending from 3DGS [KKLD23] improves results for all image quality metrics. Note that the increase in speed compared to 3DGS is mostly due to improved global as well as per-pixel culling, which reduces the number of required computations.
@@
$K=8$ 0.811 26.70 0.245 10m02s 437.9
$K=32$ 0.821 27.12 0.236 31m11s 372.4
Ours ($K=16$) 0.822 27.17 0.234 10m07s 457.3
Inference w///o Tail 0.817 26.92 0.240 10m07s 467.0
Table 4: Ablations for our hybrid transparency approach computed on the nine scenes from the Mip-NeRF360 dataset [BMV22].

### v3

@@ 1 Introduction
However, the remarkable performance of 3DGS is achieved through a series of approximations that, while effective in many scenarios, introduce limitations in multi-view consistency. One significant issue arises from the affine approximation used in projecting 3D Gaussian splats onto the image plane, causing differing pixel contributions depending on camera placement. Another one is the per-primitive sorting, which causes elliptical color patches to suddenly appear or disappear during movement (called popping [RSP24]). In order to solve the issues, it is necessary to accurately sort and evaluate Gaussians per pixel, introducing a strong bottleneck on the sorting stage. In this paper, we introduce an pipeline which improves the original 3DGS framework in eliminating these artifacts while offering superior performance through a hybrid transparency approximation.
@@
  • We evaluate our efficient implementation, which outperforms established methods significantly, while matching or exceeding their visual fidelity.
@@ 2.1 Novel View Synthesis
Alternatively, point-based models use explicit point clouds for geometry, rendering images through fast rasterization. Due to the discrete nature of point clouds, recent point-based radiance field methods employ large convolutional neural networks (CNN) for hole-filling in image-space [ASK20, RFS22, FRF23, KPLD21, HFF23] or MLPs for feature decoding [KLR22], but these NNs are computationally expensive. Addressing this limitation, Franke et al. showed that assigning a radius to each point in combination with trilinear interpolation into an image pyramid allows using a smaller CNN [FRFS24]. While enabling fast rendering, image quality of point-based models is often lacking behind implicit NeRF- and grid-based models. Very recently, Hahlbohm et al. showed that point-based models and grid-based models can be combined to improve image quality and robustness [HFK25]. Remaining issues with point-based models are temporal instabilities and the reliance on dedicated GPUs, which are both due to the used CNN. Another approach involves converting a trained NeRF model into a faster representation for inference [HSM21, YLT21]. Recent methods achieve impressive frame rates with only a small quality loss via triangle mesh baking [YHR23, RGS24] or distilling of state-of-the-art NeRF models into memory-efficient, hybrid representations [DHR24]. While they render exceptionally fast even on low-end mobile devices, a major disadvantage is their reliance on the NeRF model that needs to be trained before the – often equally expensive – baking process. Other approaches that proved effective directly optimize a triangle mesh [CFHT23] or a set of tetrahedra [KS23].
@@ 2.2 Gaussian Splatting
More recently, point-based rendering regained momentum as a powerful primitive for differentiable rendering in novel view synthesis [KPLD21, ASK20, RFS22]. In 2023, Kerbl et al. [KKLD23] re-introduced rasterization-based rendering of point primitives with anisotropic Gaussian extent in their seminal work 3D Gaussian Splatting (3DGS). Utilizing fast, tile-based rasterization [LZ21], 3DGS optimizes a set of explicit 3D Gaussians via gradient descent and an adaptive density control mechanism. Due exceptionally fast rendering during inference, 3DGS has inspired a plethora of follow-up research addressing, e.g., anti-aliasing [YCH24], compression [BKL24], dynamic scenes [LKLR24, WYF24], large scenes [KMK24], and the reliance on an initial point cloud [KRS24, NMR25]. Especially relevant to this work is StopThePop by Radl et al. [RSP24], which employs a hierarchical sorting approach to reduce popping artifacts originating from approximate sorting used in 3DGS. Notably, the per-pixel ordering resulting from this hierarchical sorting approach is not guaranteed to be fully correct. We also present a method for avoiding these popping artifacts, but achieve this by changing the blending procedure to a hybrid transparency-based approach so that our method only requires partial depth-ordering of the initial $K$ splats contributing to each pixel color. Thus, we avoid the need for a sophisticated sorting implementation which, even when compared to the approximate solution used by 3DGS, improves performance. Equally important in the context of this work is the perspectively accurate 2D Gaussian Splatting (2DGS) by Huang et al. [HYC24] as well as concurrent work that introduces ray tracing into the 3DGS framework [YSG24, MLMP24]. While the latter also provide a perspectively accurate rendering formulation by using ray tracing, they introduce a significant slowdown during both the optimization and subsequent inference. Furthermore, their evaluation of the 3D Gaussian primitives relies on matrix inversion, which is prone to numerical instabilities when Gaussians fall flat along one or more principal axes. The perspectively correct approach by Huang et al. [HYC24] circumvents these numerical instabilities by utilizing an optimized approach for 2D, i.e., degenerate 3D Gaussians [SWBG06, WHA07]. Regarding the perspectively correct rendering of non-degenerate 3D Gaussians, projecting each Gaussian onto a tangential plane of the unit sphere around the camera [HBG24] works well but significantly slows down optimization and subsequent inference. In this work, we present an approach that enables perspectively correct rendering of 3D Gaussians through ray-splat intersection: we extend the perspectively correct approach of Weyrich et al. [WHA07] to work with non-degenerate 3D Gaussians, which enables the use of ray-splat intersection for rasterization-based rendering with negligible overhead and no issues with respect to numerical instabilities.
@@ 3.2 Accurate Splat Bounding and Evaluation
In 3DGS, Kerbl et al. [KKLD23] project a 3D Gaussian onto the image plane using the Jacobian of the affine approximation of the projective transformation [ZPvBG01a]. The result is a 2D Gaussian that can easily be evaluated for each pixel and also allows for easy construction of an axis-aligned bounding box around the center (e.g., at $3\sigma$) in screen space for accelerated computation. However, this approach is not fully perspective-accurate, meaning it introduces artifacts when splats are viewed from certain angles (see Fig. 1).
@@
Next, we seek to compute the value of the 3D Gaussian for a given viewing ray. In computer graphics, ray intersections with deformed primitives, such as our 3D splats, are an established practice since the dawn of ray tracing, generally following the approach of inversely transforming a viewing ray into the undeformed space. In our case, re-using quantities of the bounding-box setup above, this would mean transforming the ray by $(VPMT)^{-1}$ into splat space, where the distance of the transformed ray to the origin yields $\rho$, Eq. 1. In order to avoid the matrix inversion that we successfully sidestepped during bounding-box calculation, we once again follow Weyrich et al. and represent a viewing ray through a pixel at $(x_{\text{s}},y_{\text{s}})$ as two perpendicular planes $\boldsymbol{\pi}_{x}=(1,0,0,-x_{\text{s}})^{\!\top\!}$ and $\boldsymbol{\pi}_{y}=(0,1,0,-y_{\text{s}})^{\!\top\!}$ in screen space. Note that we follow the common convention of representing planes as homogeneous vectors $\boldsymbol{p}=(a,b,c,d)^{\!\top}$, so that a point $\boldsymbol{x}$ lies on the plane if and only if $\boldsymbol{p}^{\!\top}\boldsymbol{x}=0$. In contrast to 3D points, transforming those planes by $(VPMT)^{-1}$ into planes $\boldsymbol{\pi}^{\text{s}}_{x/y}$ in splat space requires the inverse-transposed mapping, that is:
@@ 4 Experiments
Zip-NeRF [BMV23] 0.828 28.56 0.219 5h 5s N///a 0.878 26.75 0.233 5h 10s N///a
3DGS [KKLD23] 0.814 27.20 0.254 18m40s 3.2ms 3.31M 0.866 25.27 0.276 20m18s 4.6ms 1.50M
StopThePop [RSP24] 0.814 27.30 0.252 20m11s 4.7ms 3.28M 0.866 24.99 0.270 22m03s 4.8ms 1.49M
$\hookrightarrow$ w/// opacity decay 0.809 27.08 0.267 12m02s 2.5ms 1.71M 0.862 24.98 0.284 14m22s 3.0ms 0.81M
2DGS [HYC24] 0.795 26.81 0.297 21m08s 7.4ms 2.01M 0.851 24.55 0.314 28m45s 10.4ms 0.95M
Huang et al. [HBG24] 0.814 27.41 0.257 1h01m 10.7ms 3.30M 0.866 25.19 0.276 1h13m 15.7ms 1.49M
GOF [YSG24] 0.821 27.27 0.238 1h21m 60.0ms 2.91M 0.866 25.01 0.274 1h35m 81.3ms 1.26M
Ours 0.822 27.17 0.234 10m07s 2.2ms 2.19M 0.866 24.62 0.272 8m33s 2.6ms 0.81M
@@ 4.1 Setup
Following recent work, we evaluate on the established Mip-NeRF360 [BMV22] and Tanks and Temples [KPZK17] datasets, which provide a diverse set of 17 real scenes with varying challenges. We use all nine scenes from the Mip-NeRF360 dataset as well as all eight scenes from the intermediate set of the Tanks and Temples dataset resulting in a total of 17 scenes. This allows us to evaluate our method on a broad spectrum of challenges regarding both geometric and photometric aspects. As is common practice, we use the 4×/ 2×4$\times$/\ 2$\times$4 × / 2 × downscaled images for the outdoor///indoor scenes of the Mip-NeRF360 dataset. For the Tanks and Temples scenes, we use the full-resolution images to evaluate performance at full HD resolution. To compute meaningful quality metrics, we use the established 7:1 train///test split [BMV22] for all scenes. Central to our evaluation is the comparison against the original 3D Gaussian Splatting [KKLD23]. Furthermore, we compare against the recent StopThePop [RSP24], 2D Gaussian Splatting [HYC24], and Huang et al.’s approach [HBG24], as they also provide solutions for the problems addressed in this work. For StopThePop, we report results of both model variants the authors propose in their paper [RSP24]. We also compare against Gaussian Opacity Fields (GOF) [YSG24], a very recent work that uses ray tracing-based volume rendering. For all baselines, we use the official implementation to compute the results for all 17 scenes and use the same script for all quality metric computations. We use a single RTX 4090 for all experiments and measure frame rate by rendering the test-set images of each scene 100 times and averaging over the results [DHR24].
@@ 4.2 Implementation and Optimization
We implement our method from scratch in PyTorch and CUDA using the implementation of Kerbl et al. [KKLD23] as a reference. While the optimization pipeline and hyperparameters are mostly the same as in 3DGS, we make multiple modifications to solve technical challenges arising from our changes to the rendering pipeline. First, we tackle aliasing problems by employing our own fast CUDA implementation of the 3D filter proposed by Yu et al. [YCH24] as our inversion-free splat evaluation prevents use of the screen-space low-pass filter in EWA splatting [ZPvBG01a]. As the screen-space gradients 3DGS uses for densification are not available with our splat evaluation scheme, we follow Moenne-Loccoz et al. and instead directly use the gradient of the splat means scaled by half the respective distance to the camera in view space [MLMP24]. To prevent over-densification, we replace the opacity reset with a decay strategy [RSP24] that multiplies splat opacity by $\lambda_{o}=0.9995$ every $50$ iterations during densification. Furthermore, we adapt the densification changes from GOF to our approach [YSG24] and apply visibility score-based pruning as proposed by Niemeyer et al. [NMR25]. As we will show in our experiments, these changes are necessary to adjust 3DGS’s densification strategy to work with our ray-splat intersection-based approach. For our blending formulation, we use a core with $K=16$ and employ a second threshold $\tau_{K}=0.05$ specifying the minimum $\alpha$ of a fragment to be considered for the core. As the insertion sort that we employ for keeping track of per-pixel core causes significant divergence, we reduce the tile size to $8\times 8$ in our implementation. Lastly, we make use of an optimized SSIM implementation [MGK24] to speed up loss computation.
@@ 4.3 Results
The quantitative results in Tab. 1 show that our approach provides significantly faster training and rendering. Compared to 3DGS [KKLD23], we achieve a training speedup of 2.1×2.1$\times$2.1 × and render 1.6×1.6$\times$1.6 × faster on average. The second fastest approach is StopThePop [RSP24] with opacity decay enabled, which benefits from a large reduction in the number of splats. In terms of image quality, we observe that our method achieves about equal results across the board. However, all 3DGS-based approaches are clearly outperformed by Zip-NeRF [BMV23]. Taking our contributions into consideration, we observe that our approach outperforms approaches with perspectively correct splat evaluation. Compared to 2DGS [HYC24], we achieve better results across the board reassuring that 3D Gaussian splats are more expressive than the 2D Gaussian surfels used by 2DGS. Huang et al.’s projection method [HBG24] significantly slows down training and rendering while achieving image quality similar to that of 3DGS. GOF [YSG24] achieves slightly higher image quality but requires longer training and fails to render in real-time. It should be noted that 2DGS and GOF were fine-tuned for mesh extraction after optimization, which is an interesting topic but not a focus of this work. Seeing that we achieve similar image quality but faster training and rendering compared to either configuration of StopThePop [RSP24], the results confirm that our approach is a valid alternative to prevent popping artifacts for rasterization-based 3D Gaussian splat rendering.
3DGS [KKLD23] Huang et al. [HBG24] StopThePop [RSP24] 2DGS [HYC24] Ours Ground Truth
@@
Method Preprocess Tiling Blending Total #Splats
Bicycle [BMV22]
3DGS [KKLD23] 1.099 3.494 3.422 8.015 6.03M
StopThePop [RSP24] 1.364 1.358 5.116 7.838 6.05M
$\hookrightarrow$ w/// opacity decay 0.554 0.547 2.442 3.543 2.91M
2DGS [HYC24] 0.674 1.909 5.267 7.850 3.99M
Ours 0.556 0.754 2.007 3.317 4.35M
Francis [KPZK17]
3DGS [KKLD23] 0.235 2.078 1.644 3.957 0.81M
StopThePop [RSP24] 0.268 0.423 2.869 3.560 0.82M
$\hookrightarrow$ w/// opacity decay 0.154 0.310 1.980 2.444 0.45M
2DGS [HYC24] 0.177 4.397 8.700 13.27 0.51M
Ours 0.107 0.678 1.228 2.013 0.44M
Table 2: Breakdown of render timings in milliseconds at a resolution of 1920×1080 pixels measured on an RTX 4090 GPU. We select the scenes where the trained 3DGS model contains the highest (Bicycle) and lowest (Francis) number of splats across all tested scenes. For each method, we repeatedly render all test-set views and average measurements obtained via NVIDIA Nsight Systems across 100 runs.
Perspectively Correct Splat Evaluation. As previously discussed, our contributions can be integrated into 3D Gaussian splatting implementations independently. Our hybrid transparency approach (Sec. 3.3) is mainly a method for addressing popping artifacts without sacrificing computational efficiency. However, our perspectively accurate splat evaluation (Sec. 3.2) should prove beneficial across a much broader range of applications due to its accuracy. As can be seen in Tab. 3, using it alongside the approximately depth-ordered alpha blending from 3DGS [KKLD23] improves results for all image quality metrics. Note that a major portion of the speed increase over 3DGS comes from our approach providing tight screen-space bounds for each splat. These significantly reduce the amount of global memory accesses, sorting overhead, and the number of splats processed during blending.
@@
3DGS 0.730 24.64 0.265 21m17s 3.41ms
Ours 0.742 24.62 0.237 12m09s 2.08ms
Ours w/o per-pixel HT 0.748 24.81 0.230 17m26s 2.23ms
Table 3: We showcase the potential of our perspectively correct splat evaluation by omitting our hybrid transparency-based blending in favor of the approximately depth-ordered alpha blending from 3DGS [KKLD23]. As it forgoes per-pixel ordering, this version of our model is subject to popping artifacts, and it is slightly less efficient than our default configuration, but outperforms 3DGS across all metrics. The reported values are averaged results for the five outdoor scenes from the Mip-NeRF360 dataset [BMV22].
Hybrid Transparency Ablations. To analyze the behavior of our hybrid transparency-based blending, we perform ablations on the nine scenes from the Mip-NeRF360 dataset [BMV22]. The quantitative results shown in Tab. 4 clearly indicate that $K=16$ is the optimal choice for our approach. Note that rendering is slightly slower with $K=8$ as the number of splats created during optimization increases as core size decreases.
We also show visual comparisons in Fig. 5 showing that blending without any sorting is not a valid option as it causes the foreground to become transparent. However, we do find that background reconstruction seems to improve in this case. With our default configuration of $K=16$, omitting tail computation during inference causes a small drop in image quality, which is most noticeable in the sky. Importantly, not using the tail during training causes catastrophic failure.
@@
Hybrid Transparency
$K=8$ 0.811 26.70 0.245 10m02s 2.28ms 2.45M
$K=32$ 0.821 27.12 0.236 31m11s 2.67ms 2.10M
Ours ($K=16$) 0.822 27.17 0.234 10m07s 2.19ms 2.19M
Inference w///o tail 0.817 26.92 0.240 10m07s 2.14ms 2.19M
Densification
No 3D filter 0.777 25.58 0.303 1m44s 2.24ms 3.65M
No opacity decay 0.784 25.92 0.292 2m42s 3.81ms 4.35M
No scaled grads 0.784 26.30 0.317 4m05s 1.35ms 0.30M
No vis. pruning 0.820 26.96 0.235 12m17s 2.34ms 2.42M
Ours-7K 0.780 25.72 0.299 1m45s 2.10ms 2.09M
Ours 0.822 27.17 0.234 10m07s 2.19ms 2.19M
Table 4: Ablations for our hybrid transparency approach and our adaptation of the densification strategy from 3DGS [KKLD23] computed on the nine scenes from the Mip-NeRF360 dataset [BMV22]. Configurations marked with $\dagger$ are only trained for $7,000$ iterations as the optimization would otherwise run out of memory due to over-densification.
Densification Ablations. As described in Sec. 4.2, we make multiple adjustments to the adaptive density control mechanism proposed by Kerbl et al. [KKLD23]. The ablations in Tab. 4 clearly show why this is necessary for our ray-splat intersection-based approach. Not scaling the gradient of the splat means by half the respective distance to the camera [MLMP24] results in significant under-reconstruction, whereas dropping either the 3D filter [YCH24] or using the standard opacity reset [KKLD23] every 3,000 iterations instead of an opacity decay [RSP24] results in the optimization running out of memory due to over-reconstruction. Furthermore, the visibility score-based pruning [NMR25] simultaneously improves quality and speed as it helps to prevents over-reconstruction by pruning low-opacity splats, which would otherwise cause overfitting.
5 Discussion